# Jiafei Duan 1    Wentao Yuan 1∗    Wilbert Pumacay 2    Yi Ru Wang 1 Kiana Ehsani 3    Dieter Fox 1,4   Ranjay Krishna 1,3

### v2

@@ 4.2 Behavior cloning with demonstrations from Manipulate-Anything
Generated data Put_block Play_jenga Open_jar Close_box Open_box Pickup_cup
VoxPoser[3] 2.67±$\pm 2.31$ $\pm 4.00$
CAP[4] 6.67±$\pm 2.31$ 14.67±$\pm 12.86$
MA (Ours) 85.33±$\pm 10.07$ 81.33±$\pm 2.31$ 21.33±$\pm 10.07$ 42.67±$\pm 8.33$ 30.67±$\pm 11.55$ 54.00±$\pm 12.49$
RLBench[33] 20.00±$\pm 18.33$ 81.33±$\pm 9.24$ 58.67±$\pm 45.49$ 68.00±$\pm 24.98$ 14.67±$\pm 6.11$ 54.67±$\pm 23.09$
Generated data Take_umbrella Sort_mustard Open_wine Lamp_on Put_knife Pick_&_lift
VoxPoser[3] $\pm 4.00$ $\pm 0.00$ 1.33±$\pm 2.31$ 5.33±$\pm 4.62$ 1.33±$\pm 2.31$ 5.67±$\pm 1.64$
CAP[4] 13.33±$\pm 10.06$ 8.00±$\pm 16.00$ 9.33±$\pm 6.11$ 46.67±$\pm 2.31$
MA (Ours) 84.00±$\pm 6.93$ 53.33±$\pm 6.11$ 86.67±$\pm 6.11$ 89.33±$\pm 6.11$ 8.00±$\pm 4.00$ 33.33±$\pm 2.31$
RLBench[33] 58.67±$\pm 50.80$ 53.33±$\pm 34.02$ 86.67±$\pm 12.86$ 84.00±$\pm 13.86$ 30.67±$\pm 10.07$ 62.67±$\pm 9.24$
@@
Training and evaluation protocol. Using the generated demonstrations, we train a Perceiver-Actor (PerAct) model, which is a transformer-based robotic manipulation behavior cloning model [34]. The model expects tokenized voxel grids and language instructions as inputs and predicts discretized voxel grid 6 DoF poses and gripper states. For all the generated training datasets, we train a multi-task PerAct policy with a batch size of $4$ for $30$k iterations. To ensure consistent evaluation, we generate one set of testing environments with RLBench. We evaluate the last checkpoint from each of the trained policies. Each policy is evaluated for $25$ episodes across each task using $3$ different seeds. We measure the success rate based on the simulation-defined success condition.
Results: Policies trained using Manipulate-Anything data perform similarily to policies trained using human demonstrations ($p=0.973$) (Table 2). Training on either Manipulate-Anything or on human demonstrations results in a performance difference of a mere $0.27$% across all tasks. Furthermore, models trained on data from the baselines exhibit a statistically lower performance ($p\leq 0.01$ for both VoxPoser and CAP). One of the main factors potentially contributing to the differences in the performance could be that Manipulate-Anything generates diverse expert trajectories that are preferable to humans. This can be seen in Fig. 5, which shows the action distribution of the generated data by different methods for the same given tasks. Additionally, our generated data recorded the lowest Chamfer Distance (CD) of 0.056 with human-generated demonstrations data. We also observed that the policy trained on MA data achieves a lower standard deviation of $3.39$ across all tasks compared to zero-shot performance of $8.48$. This suggests the benefits of training over generated data instead of relying solely on zero-shot deployment.
@@ 4.3 Real-world experiments
Environment and tasks. We employ a Franka Panda manipulator equipped with a parallel gripper. We use a front-facing Kinect 2 RGB-D camera. To generate multi-view inputs for the Manipulate-Anything framework, we re-render virtual viewpoints from the generated point cloud, similar to prior work [51]. We select 5 representative real world tasks: open_jar, sort_objects, correct_dices, open_drawer, and on_lamp, all conditioned on language instructions. We evaluate each task for $10$ episodes, with varying object poses across $3$ trials of evaluation.
@@
Open_drawer Sort_object On_lamp Open_jar Correct_dice
CAP (0-shot) $00.00\pm 0$ 0.00 $13.33\pm 5.77$ $00.00\pm 0$ 0.00 $6.67\pm 5.77$ $6.67\pm 5.77$
MA (0-shot) 36.67±$\pm 5.77$ 60.00±$\pm 10.00$ 26.67±$\pm 11.55$ 40.00±$\pm 10.00$ 53.33±$\pm 5.77$
PerAct (MA data) $50.00\pm 0.00$ $33.33\pm 5.77$ $46.67\pm 5.77$ $56.67\pm 5.77$ $60.00\pm 0.00$
PerAct (Human data) 53.33±$\pm 11.55$ 36.67±$\pm 5.77$ 60.00±$\pm 0.00$ 76.67±$\pm 5.77$ 80.00±$\pm 10.00$
Results: Manipulate-Anything is able to generate successful demonstrations for each of the $5$ real world tasks. Even for the worst-performing task, Manipulate-Anything achieves a success rate of more than $25\%$. Our approach outperforms CAP by 38%. Consistent with the simulation results, training with the data generated by Manipulate-Anything produces a more robust policy compared to performing zero-shot. Additionally, in 4 out of 5 tasks, the trained policies perform better than the zero-shot approach. The policy underperforms on the sort_object task, because it requires longer-horizon memory —a known limitation pointed out in PerAct [34].
@@ 4.4 Ablations
For effective real-world deployment of Manipulate-Anything, it’s crucial that the collected data supports scaling of robotics transformers and offers diverse skills and interacted objects. We conducted an ablation study to evaluate the quality of Manipulate-Anything-generated data for scaling and its generalization to language instruction changes. For scaling, we generated behavior cloning data, ranging from 1 to 100 training demonstrations from RLBench and Manipulate-Anything for a single task, and trained a PerAct policy. For generalization, we varied the sort_mustard task with different language instructions and target objects. We compared our approach to VoxPoser to assess robustness to object and language instruction changes. Further implementation details are in the supplementary materials. Result: Our scaling experiments demonstrate that generating more training data via Manipulate-Anything improves PerAct policy performance (Fig. 6). The data from our approach shows a better rate of change with a slope of 0.503 for a linear fit, compared to 0.197 for RLBench-generated data. Additionally, Manipulate-Anything data is more generalizable and robust to language instruction changes, outperforming VoxPoser in task success across language and object variations. Detailed results in the appendix.
@@ 5 Discussion
Limitations. Manipulate-Anything relies on the availability of VLMs. While this can pose a dependence on the foundational models with the rise of the open VLMs, we believe this issue will be addressed soon. Future work. With the enhancements of large foundational models, Manipulate-Anything, due to its modularity, will continue to grow and scale up to more complex tasks. Conclusion. Manipulate-Anything is a scalable environment-agnostic approach for generating 0-shot demonstration for robotic tasks without the use of privileged environment information. Manipulate-Anything uses VLMs to do high level planning and scene understanding and is capable of error recovery. This enables high quality data generation for behavior cloning that can achieve better performance that using human data.
6 Acknowledgement
Jiafei Duan is supported by the Agency for Science, Technology and Research (A*STAR) National Science Fellowship. Wilbert Pumacay is supported by grant 234-2015-FONDECYT from Cienciactiva of the National Council for Science, Technology and Technological Innovation (CONCYTEC-PERU). This project is partially supported by Amazon Science. We would also like to thank Winson Han from AI2 for helping with the figure and icon design.

### v3

@@ 4.2 Behavior cloning with demonstrations from Manipulate-Anything
Data Models Put_block Play_jenga Open_jar Close_box Open_box Pickup_cup
VoxPoser[3] PerAct[34] 2.67±$\pm 2.31$ $\pm 4.00$
CAP[5] PerAct[34] 6.67±$\pm 2.31$ 14.67±$\pm 12.86$
Scaling-up [4] PerAct[34] 22.67±$\pm 15.14$ 5.33±$\pm 6.11$ 14.67±$\pm 2.31$
MA (Ours) PerAct[34] 85.33±$\pm 10.07$ 81.33±$\pm 2.31$ 21.33±$\pm 10.07$ 42.67±$\pm 8.33$ 30.67±$\pm 11.55$ 54.00±$\pm 12.49$
RLBench[33] PerAct[34] 20.00±$\pm 18.33$ 81.33±$\pm 9.24$ 58.67±$\pm 45.49$ 68.00±$\pm 24.98$ 14.67±$\pm 6.11$ 54.67±$\pm 23.09$
VoxPoser[3] RVT-2[59] 73.33±$\pm 2.31$ 2.67±$\pm 2.31$
CAP[5] RVT-2[59] 78.66±$\pm 8.32$ 77.33±$\pm 19.73$
Scaling-up [4] RVT-2[59] 38.67±$\pm 2.31$ 33.33±$\pm 2.31$ 92.00±$\pm 4.00$
MA (Ours) RVT-2[59] 85.33±$\pm 2.31$ 82.67±$\pm 2.31$ 78.67±$\pm 10.06$ 82.67±$\pm 2.31$ 24.00±$\pm 4.00$ 97.33±$\pm 2.31$
RLBench[33] RVT-2[59] 86.67±$\pm 2.31$ 85.33±$\pm 2.31$ 81.33±$\pm 6.11$ 76.00±$\pm 4.00$ $\pm 4.00$ 97.33±$\pm 2.31$
Data Models Take_umbrella Sort_mustard Open_wine Lamp_on Put_knife Pick_&_lift
VoxPoser[3] PerAct[34] $\pm 4.00$ $\pm 0.00$ 1.33±$\pm 2.31$ 5.33±$\pm 4.62$ 1.33±$\pm 2.31$ 5.67±$\pm 1.64$
CAP[5] PerAct[34] 13.33±$\pm 10.06$ 8.00±$\pm 16.00$ 9.33±$\pm 6.11$ 46.67±$\pm 2.31$
Scaling-up [4] PerAct[34] $\pm 4.00$ $\pm 0.00$ 81.33±$\pm 12.86$ 76.00±$\pm 4.00$ 5.33±$\pm 2.31$ 53.33±$\pm 10.06$
MA (Ours) PerAct[34] 84.00±$\pm 6.93$ 53.33±$\pm 6.11$ 86.67±$\pm 6.11$ 89.33±$\pm 6.11$ 8.00±$\pm 4.00$ 33.33±$\pm 2.31$
RLBench[33] PerAct[34] 58.67±$\pm 50.80$ 53.33±$\pm 34.02$ 86.67±$\pm 12.86$ 84.00±$\pm 13.86$ 30.67±$\pm 10.07$ 62.67±$\pm 9.24$
VoxPoser[3] RVT-2[59] 5.33±$\pm 6.11$ 1.33±$\pm 2.31$ 1.33±$\pm 2.31$ 2.67±$\pm 2.31$ 1.33±$\pm 2.31$ 17.33±$\pm 2.31$
CAP[5] RVT-2[59] 89.33±$\pm 6.11$ 85.33±$\pm 8.32$ 52.00±$\pm 10.58$ 82.66±$\pm 20.53$
Scaling-up [4] RVT-2[59] 94.67±$\pm 4.62$ 24.00±$\pm 4.00$ 62.67±$\pm 2.31$ $\pm 2.31$ 53.33±$\pm 2.31$ 80.00±$\pm 6.93$
MA (Ours) RVT-2[59] 94.67±$\pm 2.31$ 73.33±$\pm 2.31$ 93.33±$\pm 6.11$ 84.00±$\pm 10.58$ 69.33±$\pm 12.85$ 82.67±$\pm 12.22$
RLBench[33] RVT-2[59] 97.33±$\pm 2.31$ 69.33±$\pm 8.33$ 88.00±$\pm 8.00$ 93.33±$\pm 4.62$ 72.00±$\pm 10.58$ 64.00±$\pm 10.58$
@@
Training and evaluation protocol. We train two models using the generated demonstrations: the Perceiver-Actor (PerAct) model [34], a transformer-based robotic manipulation behavior cloning model that expects tokenized voxel grids and language instructions as inputs and predicts discretized voxel grid 6 DoF poses and gripper states, and also RVT-2 model [59], a multi-view transformer-based BC model. The RVT-2 model uses tokenized image patches and CLIP-encoded language instructions as input to predict keypoint actions as translation heatmaps, discretized rotation in Euler angles, and the gripper’s binary state. Notably, RVT-2 is currently the highest-performing model on the RLBench benchmark. For all the generated training datasets, we train a multi-task PerAct policy with a batch size of $4$ for $30$k iterations on a single RTX A100, and RVT-2 with a batch size of 24 for $3.3$k iterations on 4 A100s. To ensure consistent evaluation, we generate one set of testing environments with RLBench. We evaluate the last checkpoint from each of the trained policies. Each policy is evaluated for $25$ episodes across each task using $3$ different seeds. We measure the success rate based on the simulation-defined success condition.
Results: Policies trained using Manipulate-Anything data perform similarly to policies trained using hand-scripted demonstrations ($p=0.973$) for PerAct (Table 2). Training on either Manipulate-Anything or on hand-scripted demonstrations results in a performance difference of a mere $0.27$% across all tasks. Furthermore, models trained on data from the baselines exhibit a statistically lower performance ($p\leq 0.01$ for VoxPoser, Scaling-up and CAP). One of the main factors potentially contributing to the differences in the performance could be that Manipulate-Anything generates diverse expert trajectories that are preferable to humans. This can be seen in Fig. 5, which shows the action distribution of the generated data by different methods for the same given tasks. Additionally, our generated data recorded the lowest Chamfer Distance (CD) of 0.056 with human-generated demonstrations data. We also observed that the policy trained on MA data achieves a lower standard deviation of $3.39$ across all tasks, compared to the zero-shot performance standard deviation of $8.48$. This suggests the benefits of training over generated data instead of relying solely on zero-shot deployment.
@@ 4.3 Real-world experiments
Environment and tasks. We employ a Franka Panda manipulator equipped with a parallel gripper. We use a front-facing Kinect 2 RGB-D camera. To generate multi-view inputs for the Manipulate-Anything framework, we re-render virtual viewpoints from the generated point cloud, similar to prior work [53]. We selected 7 representative real-world tasks, both prehensile and non-prehensile: open_jar, sort_objects, correct_dices, open_drawer, on_lamp, press_switch, and close_laptop, all conditioned on language instructions. Each task was evaluated over 10 episodes with varying object poses across 3 trials.
@@
Open_drawer Sort_object On_lamp Open_jar Correct_dice Press_switch Close_laptop
CAP (0-shot) $00.00\pm 0$ 0.00 $13.33\pm 5.77$ $00.00\pm 0$ 0.00 $6.67\pm 5.77$ $6.67\pm 5.77$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
MA (0-shot) 36.67±$\pm 5.77$ 60.00±$\pm 10.00$ 26.67±$\pm 11.55$ 40.00±$\pm 10.00$ 53.33±$\pm 5.77$ 20.00±$\pm 10.00$ 33.33±$\pm 5.77$
PerAct (MA data) $50.00\pm 0.00$ $33.33\pm 5.77$ $50.00\pm 0.00$ $56.67\pm 5.77$ $60.00\pm 0.00$ $56.67\pm 5.77$ $33.33\pm 5.77$
PerAct (Human data) 53.33±$\pm 11.55$ 36.67±$\pm 5.77$ 60.00±$\pm 0.00$ 76.67±$\pm 5.77$ 80.00±$\pm 10.00$ 33.33±$\pm 5.77$ 53.33±$\pm 5.77$
Results: Manipulate-Anything is able to generate successful demonstrations for each of the $7$ real world tasks. Even for the worst-performing task, Manipulate-Anything achieves a success rate of more than $25\%$. Our approach outperforms CAP by 38%. Consistent with the simulation results, training with the data generated by Manipulate-Anything produces a more robust policy compared to performing zero-shot. Additionally, in 4 out of 5 tasks, the trained policies perform better than the zero-shot approach. The policy under-performs on the sort_object task, because it requires longer-horizon memory —a known limitation pointed out in PerAct [34].
Figure 6: Error breakdown. Error breakdown for three tasks from simulation.
@@ 4.4 Ablations
For effective real-world deployment of Manipulate-Anything, it’s crucial that the collected data supports scaling of robotics transformers and offers diverse skills and interacted objects. We conducted an ablation study to evaluate the quality of Manipulate-Anything-generated data for scaling and its generalization to language instruction changes. For scaling, we generated behavior cloning data, ranging from 1 to 100 training demonstrations from RLBench and Manipulate-Anything for a single task (put_block), and trained a PerAct policy. For generalization, we varied the sort_mustard task with different language instructions and target objects. We compared our approach to VoxPoser to assess robustness to object and language instruction changes. Further implementation details are in the supplementary materials. Result: Our scaling experiments demonstrate that generating more training data via Manipulate-Anything improves PerAct policy performance (Fig. 6). The data from our approach shows a better rate of change with a slope of 0.503 for a linear fit, compared to 0.197 for RLBench-generated data. Additionally, Manipulate-Anything data is more generalizable and robust to language instruction changes, outperforming VoxPoser in task success across language and object variations. Detailed results in the appendix.
4.5 Error breakdown
We examine the potential errors introduced by each respective VLM and explore ways to further improve the overall system. We conducted simulation experiments where tasks could be easily reset, focusing on two major components where VLMs make decisions: "Perception error" and "Reasoning error." "Perception error" refers to mistakes made by the VLM in detecting target objects and selecting optimal viewpoints for generating task-specific grasp pose, primarily related to the selection of the Multi-viewpoint VLM. "Reasoning error" involves the Sub-task Verification Module, where the VLM is responsible for deciding if sub-tasks have been successfully completed. Due to resource constraints, we conducted experiments on the play_jenga task. We systematically replaced the VLMs in each component with a human, allowing the human to make decisions. By comparing the system’s performance with human decision-making to that with VLMs, we were able to quantify the errors caused by the VLMs.
@@ 5 Discussion
Limitations. Despite the promising results, Manipulate-Anything has several limitations. First, Manipulate-Anything depends on the availability of large language models (LLMs), which introduces reliance on foundational models. However, with the ongoing development of open-source LLMs, this issue is likely to diminish over time. Second, Manipulate-Anything struggles with dynamic manipulation tasks and non-prehensile tasks, as it cannot generate alternative grasp poses for these scenarios. Third, the system’s highly modular nature, integrating multiple VLMs, can lead to compounding errors when generating zero-shot trajectories. Emerging specialized VLMs [51] and others which may help address this issue. Finally, manual prompt engineering for in-context learning is still required. Nonetheless, recent advancements in alignment and prompting techniques [60, 61, 62] offer potential solutions to reduce the effort involved in prompt engineering.
Conclusion. Manipulate-Anything is a scalable environment-agnostic approach for generating zero-shot demonstration for robotic tasks without the use of privileged environment information. Manipulate-Anything uses LLMs to do high level planning and scene understanding and is capable of error recovery. This enables high quality data generation for behavior cloning that can achieve better performance that using human data.